Comment epi2 SNF 2008-10-06 09:42:13: Update overtemp

eenriquez at eenriquez at
Mon Oct 6 09:42:13 PDT 2008

Reset the lamp housing over temperature safety switch.  Found the dome is heavily encrusted with deposition.   Unable to run an HCl clean since the dome now is opaque.  Will purge the chamber for 24 hours before opening.  Will assess the damage to the chamber components tomorrow.  

More information about the epi2-pcs mailing list